OpenStudy (anonymous):

find the next three terms in the pattern 9,4,-1,-6,-11,.... then describe the pattern.

OpenStudy (mathmath333):

its Arithmatic Series with common difference 5

OpenStudy (mathmath333):

9-5=4 4-5=-1 -1-5=-6 -6-5=-11 -11-5=-16